Find out how long it takes a $3500 investment to double of it is invested at 8% compunded semiannuakky. Round to the nearest ten of a year.

Answer by jorel1380(3719)

If the money is invested at 8%,compounded semiannually, then the formula for its value after n years is:
fv=3500(1+.08/2)^2n
So:
7000=3500(1.04)^2n
2=1.04^2n
ln 2=ln 1.04^2n
ln 2=2n ln 1.04
ln 2/2 ln 1.04=n
n=8.83 years. ☺☺☺☺